The cost of debt   Gronseth Drywall​ Systems, Inc., is in discussions with its investment bankers regarding the issuance of new bonds. The investment banker has informed the firm that different maturities will carry different coupon rates and sell at different prices. The firm must choose among several alternatives. In each​ case, the bonds will have a ​$1000 par value and flotation costs will be ​$35 per bond. The company is taxed at 28​%. Use the approximation formula to calculate the ​after-tax cost of financing with the following alternative.  ​(Click on the icon located on the​ top-right corner of the data table below in order to copy its contents into a​ spreadsheet.) Coupon rate 12% Time to maturity 15 years Premium or discount $210
Answer #1

Future Value = $1,000
Present value = ($1,000 - $35) + $210 = $1,175
Nper = 15
PMT = $1,000 * 12% = $120
Before tax cost of debt can be calculated using the following excel formula:
=RATE(nper,pmt,pv,fv)
=RATE(15,-120,1175,-1000)
= 9.734%

After tax cost of debt = 9.734 * (1 - 0.28) = 7.0085% or 7%

After tax cost of debt = 7.0085% or 7%